OpenStudy (anonymous):

3 radical x - 5 radical x=?

jimthompson5910 (jim_thompson5910):

\[\Large 3\sqrt{x} - 5\sqrt{x} = (3 - 5)\sqrt{x}\] \[\Large 3\sqrt{x} - 5\sqrt{x} = (-2)\sqrt{x}\] \[\Large 3\sqrt{x} - 5\sqrt{x} = -2\sqrt{x}\]